substitution and elimination
2x+y=-5
-x+3y=6

Answer by checkley77(12844)

2x+y=-5
-x+3y=6 multiply by 2 & add.
-2x+6y=12
2x+y=-5
-----------------
7y=7
y=7/7
y=1 ans.
2x+1=-5
2x=-5-1
2x=-6
x=-6/2
x=-3 ans.
Proof:
-(-3)+3*1=6
3+3=6
6=6